# Breaking the Circle: An Autonomous Control-Switching Strategy for Stable Orographic Soaring in MAVs

### v1

@@ IV SIMULATION
We implemented the new control effectiveness model and switching method in our MATLAB simulation environment. The environment features a wind field that was precomputed using ANSYS Fluent [40], based on the geometry of the TU Delft Open Jet Facility (OJF).
@@
    BASE Controller: The original INDI-based controller from AOSoar [26] without any modifications.
@@
Figure 9: Position standard deviation (STD) over time for each control configuration. The plot shows the evolution of 3D position variability across non-diverged runs from 300 randomized reference positions. Lower STD values indicate greater soaring stability. Diverged cases were excluded to ensure a fair comparison.
@@ V-B MAV Hardware
Two platforms of different sizes and configurations were used for testing to evaluate the generality of the proposed control method. The first was the Eclipson Model C [41], shown in figure 13.(a), the same model used in our previous research [26], featuring a wingspan of $1100mm$, a wing surface area of $18dm^{2}$, and a total weight of $716g$. The second platform was the Eachine & ATOMRC Seal G1500, shown in figure 13.(b), with a wingspan of $1500mm$, a wing area of $25.7dm^{2}$, and a total weight of $1210g$. Both MAVs were equipped with an elevator, rudder, and ailerons for control, along with a single electric motor for throttle.
The electronics setup was similar across both platforms. We used the Paparazzi autopilot [42] running on a Pixhawk 4. Airspeed and angle-of-attack sensors were installed in different locations on each MAV to accommodate differences in available space and airframe design. Additionally, Opti-Track markers were attached to both MAVs for accurate indoor position tracking.
Figure 12: Ramp used to generate updraft (left) and overall test setup, including the wind section, at the TU Delft Open Jet Facility (right).
@@ VI RESULTS
In this section, we present the experimental results from the wind tunnel test. Flight results from both platforms showed that the SAOS reduced position oscillations and throttle usage, allowing the MAV to maintain a steadier soaring flight. Switching back to the BASE controller from the SAOS immediately increased oscillations and position errors, indicating its difficulty in handling control conflicts. Overall, the results confirm that the SAOS improves the circling behavior in soaring flight.
@@ VII DISCUSSION
The flight test results confirm that the proposed SAOS significantly improves both stability and energy efficiency during orographic soaring, compared to the BASE controller. By dynamically switching control authority based on throttle availability, the MAV was able to eliminate persistent circling behavior, reduce roll oscillations, and maintain its position with minimal energy expenditure. These findings highlight the importance of adapting control strategies in wind-hovering scenarios where the throttle is saturated and control surface authority is limited. The use of angle-of-attack sensing and enhanced aerodynamic modeling further contributed to stall prevention and improved force estimation. Notably, the method’s effectiveness across two MAV platforms with distinct geometries and dynamics supports its generality and potential for broader application.
@@ VIII CONCLUSION
Acknowledgments
We would like to thank Simon Cajagi for helping with the wind tunnel experiment, and MAVLab members for their support.

### v2

@@ IV SIMULATION
We implemented the new control effectiveness model and switching method in our MATLAB simulation environment. The environment features a wind field that was precomputed using ANSYS Fluent [ansysfluent], based on the geometry of the Open Jet Facility (OJF).
@@
    BASE Controller: The original INDI-based controller from AOSoar [hwang2023aosoar] without any modifications.
@@
Figure 9: Position standard deviation (STD) over time for each control configuration. The plot shows the evolution of 3D position variability across non-diverged runs from 300 randomized reference positions. Lower STD values indicate greater soaring stability. Diverged cases were excluded to ensure a fair comparison; the convergence and divergence rates for each configuration are provided in Table I.
@@ V-B MAV Hardware
Two platforms of different sizes and configurations were used for testing to evaluate the generality of the proposed control method. The first was the Eclipson Model C [eclipson], shown in figure 13 (a)., featuring a wingspan of $1100mm$, a wing surface area of $18dm^{2}$, and a total weight of $716g$. The second platform was the Eachine & ATOMRC Seal G1500, shown in figure 13 (b)., with a wingspan of $1500mm$, a wing area of $25.7dm^{2}$, and a total weight of $1210g$. Both MAVs were equipped with an elevator, rudder, and ailerons for control, along with a single electric motor for throttle.
The electronics setup was similar across both platforms. We used the Paparazzi autopilot [hattenberger2014using] running on a Pixhawk 4. Airspeed and angle-of-attack sensors were installed in different locations on each MAV to accommodate differences in available space and airframe design. Additionally, Opti-Track markers were attached to both MAVs for accurate indoor position tracking.
Figure 12: Ramp used to generate updraft (left) and overall test setup, including the wind section, at the Open Jet Facility (right).
@@ VI RESULTS
In this section, we present the experimental results from the wind tunnel test. Flight results from both platforms showed that the SAOS reduced position oscillations and throttle usage, allowing the MAV to maintain a steadier soaring flight. Switching back to the BASE controller from the SAOS immediately increased oscillations and position errors, indicating its difficulty in handling control conflicts. Overall, the results confirm that the SAOS effectively combats the circling behavior in soaring flight.
@@ VII DISCUSSION
The flight test results confirm that the proposed SAOS significantly improves both stability and energy efficiency during orographic soaring, compared to the BASE controller. By dynamically switching control authority based on throttle availability, the MAV was able to eliminate persistent circling behavior, reduce roll oscillations, and maintain its position with minimal energy expenditure. These findings highlight the importance of adapting control strategies in wind-hovering scenarios where the throttle is saturated and control surface authority is limited. While the indoor test flights and simulations did not reveal a significant advantage from angle-of-attack limiting, this stall-prevention feature is expected to be more beneficial in outdoor environments. In such settings, the MAV typically undergoes more aggressive maneuvers and experiences larger position deviations, increasing the risk of stall. Consequently, stall prevention may play a more critical role in maintaining stability during outdoor flights. Notably, the method’s effectiveness across two MAV platforms with distinct geometries and dynamics supports its generality and potential for broader application.